# Attractor Memory for Long-Term Time Series Forecasting: A Chaos Perspective

### v4

@@ 3.3 System Evolution by Attractors
  • Direct Evolution: We employ each representation $x_{t}/s_{t}\in\mathbb{R}^{D\times N}$ as the feature to partition adjacent points in the dynamic system trajectory into the same attractor using the K-means method. Subsequently, we evolve these points using a local operator $\mathcal{K}^{(i)}$.
  • Frequency-enhanced Evolution: Inspired by neuroscience [6], where attractor structures are amplified in the frequency domain, we first obtain the frequency domain representation of the dynamic structure through Fourier transformation. Considering the dominant modes as attractors, we employ ${\mathcal{K}^{(i)}}$ to drive the system’s evolution in the frequency domain.
  • Hopfield Evolution: Hopfield networks [16] are designed specifically for attractor memory retrieval (see Appendix A.2). In our approach, we utilize a modern version [33] of the Hopfield network for the evolution of dynamic systems, employing cross-attention operations. We treat the trainable attractor library as the Key and Value, while different scales of dynamic structure representations $\{s^{\theta},s^{2\theta},...,s^{2^{L}\theta},x^{2^{L}\theta}\}$ serve as the Query, enabling sequence-to-sequence evolution.
@@
Explicit Evolution. Initially, the finest dynamic representation $x^{\theta}$ is obtained using $\overline{\bm{K}}$, and then expanded to multiple scales ${s^{\theta},s^{2\theta},...,s^{2^{L}\theta},x^{2^{L}\theta}}$. By applying the Fourier Transform, we select $M$ low-frequency components as the primary modes. Each mode undergoes linear evolution $\mathcal{K}^{(i)}=W_{i}\in\mathbb{C}^{N\times N}$, followed by back projection to the original scale.
Implicit Evolution. However, explicit evolution methods inevitably increase the time complexity. To address this issue, inspired by the Blelloch algorithm and hierarchical projection, which both utilize tree-like structures, we propose an implicit evolution method.
A Blelloch scan [3, 36] (Figure 3(d)) defines a binary operator $q_{i}\bullet q_{j}:=(q_{j,a}\odot q_{i,a},~{}q_{j,a}\otimes q_{i,b}+q_{j,b})$ used to compute the linear recurrence $x_{k}=\overline{\bm{A}}x_{k-1}+\overline{\bm{B}}u_{k}$. We take $L=4$ as an example.
@@
The process commences by computing the values of variable $x$ at even positions through an upward sweep. Subsequently, these even position values are employed during a downward sweep to calculate the values at odd positions. The corresponding value of $x$ resides in the right node of $r$. Thus, we can modify the binary operator as $q_{i}\bullet q_{j}:=(q_{j,a}\odot q_{i,a},~{}H_{i}\otimes(q_{j,a}\otimes q_{i,b}+q_{j,b}))$, thereby implicitly integrating hierarchical projection into the scanning operation. This leads to the scales of $x_{i}$ is:
@@
In Attraos, we utilize this indirect and efficient hierarchical projection as the default setting. Ultimately, Attraos projects from the polynomial spectral space back to the phase space by employing another gating projection $\bm{W}_{out}\in\mathbb{R}^{B\times L\times N\times 1}$ (13) to $x\in\mathbb{R}^{B\times L\times D\times N}$, and derives the prediction results using an observation function parameterized by $\bm{W}_{h}\in\mathbb{R}^{LD\times H}$ (flattening the patches).
@@ 4 Experiments
In this section, we commence by conducting a comprehensive performance comparison of Attraos against other state-of-the-art models in seven mainstream LTSF datasets along with two typical chaotic datasets, namely Lorenz96-3d and Air-Convection, followed by ablation experiments pertaining to model architectures. Furthermore, leveraging the properties of chaotic dynamic systems, we explore their extended applications, including experiments on chaotic evolution, representation, modulation, and reconstruction (the latter three are in Appendix E). Finally we provide the complexity analysis and robustness analysis. For detailed information regarding baseline models, dataset descriptions, experimental settings, hyper-parameter analysis, please refer to Appendix D.
@@ 4.1 Overall Performance
Model Attraos Mamba4TS S-Mamba RWKV-TS GPT-TS Koopa InvTrm PatchTST DLinear (Ours) [39]) [17] [17] [27] [26] [28] [43] Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 0.423 0.420 0.444 0.438 0.459 0.453 0.454 0.446 0.457 0.450 0.450 0.443 0.463 0.454 0.434 0.435 0.462 0.458 ETTh2 0.372 0.399 0.386 0.410 0.381 0.407 0.375 0.402 0.389 0.414 0.397 0.417 0.383 0.407 0.380 0.406 0.564 0.520 ETTm1 0.382 0.391 0.396 0.406 0.399 0.407 0.391 0.403 0.396 0.401 0.395 0.403 0.407 0.412 0.403 0.398 0.403 0.406 ETTm2 0.280 0.324 0.299 0.343 0.289 0.333 0.285 0.330 0.294 0.339 0.281 0.326 0.291 0.335 0.283 0.329 0.345 0.396 Exchange 0.349 0.395 0.364 0.405 0.364 0.407 0.406 0.439 0.371 0.409 0.390 0.424 0.366 0.416 0.383 0.416 0.346 0.416 Crypto 0.187 0.157 0.193 0.162 0.198 0.163 0.190 0.159 0.196 0.164 0.199 0.165 0.196 0.164 0.192 0.161 0.201 0.176 Weather 0.246 0.271 0.258 0.280 0.252 0.277 0.256 0.280 0.279 0.279 0.247 0.273 0.260 0.280 0.258 0.280 0.267 0.319
Mainstream LTSF Datasets. As depicted in Table 1, we can observe that: (a) Attraos consistently exhibits the best performance, closely followed by RWKV-TS and Koopa, which underscores the crucial role of modeling temporal dynamics in LTSF tasks. (b) The models based on state space models (Mamba4TS, S-Mamba) generally outperform the Transformer-based models (PatchTST, InvTrm), indicating the potential superiority of state space models as fundamental frameworks for temporal modeling. (c) Performance of the GPT-TS model, which relies on a pretrained large language model, is relatively average, suggesting the inherent challenge in directly capturing the dynamics of temporal data using such models. A promising avenue for future research lies in training a comprehensive foundational temporal dynamics model on a large scale or leverage pre-training to obtain an attractor tokenizer that is better suited for inputs to the large language model.
@@
Model Attraos Mamba4TS S-Mamba RWKV-TS Koopa InvTrm PatchTST DLinear (Ours) (Temporal Emb.) [39]) [17] [27] [26] [28] [43] Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Lorenz96-3d 96 0.844 0.684 0.892 0.721 0.925 0.744 0.894 0.722 0.891 0.736 0.963 0.786 0.929 0.756 0.881 0.750 192 0.835 0.662 0.910 0.748 0.917 0.761 0.894 0.744 0.881 0.752 0.944 0.811 0.899 0.714 0.910 0.753 336 0.837 0.681 0.943 0.772 0.968 0.788 0.982 0.823 0.914 0.753 0.997 0.841 0.922 0.787 0.893 0.737 720 0.872 0.739 0.996 0.814 1.135 0.940 1.058 0.921 0.989 0.801 1.129 0.955 0.971 0.828 0.927 0.806 AVG 0.847 0.692 0.935 0.764 0.986 0.808 0.957 0.803 0.919 0.761 1.008 0.848 0.930 0.771 0.903 0.762 Air 96 0.437 0.303 0.451 0.314 0.468 0.329 0.447 0.308 0.443 0.307 0.470 0.337 0.465 0.331 0.441 0.325 192 0.455 0.321 0.472 0.331 0.481 0.340 0.467 0.328 0.451 0.329 0.485 0.349 0.477 0.341 0.460 0.338 336 0.456 0.334 0.468 0.342 0.485 0.351 0.461 0.339 0.468 0.342 0.499 0.363 0.484 0.353 0.461 0.341 720 0.466 0.355 0.492 0.379 0.501 0.386 0.482 0.367 0.488 0.369 0.516 0.401 0.504 0.392 0.474 0.359 AVG 0.454 0.328 0.471 0.342 0.484 0.352 0.464 0.336 0.463 0.337 0.493 0.363 0.483 0.354 0.459 0.341
@@ 4.2 Further Analysis
Ablation Studies.  Next, we turn off each module of Attraos to assess their individual effects. As shown in Table 3, we observe consistent performance decline of Attraos when deleting each module: (a) The removal of Phase Space Reconstruction exhibits the most severe performance degradation, indicating that the process of reconstructing the dynamical structure through PSR forms the foundation for Attraos’ efficient capture of attractor structures. (b) Multi-scale hierarchical projection effectively captures the complex topological structure of the singular attractor, leading to improved performance. However, overfitting may occur in certain prediction lengths. (c) Time-varying $\bm{B}$ and $\bm{W}_{out}$ acts as a gating attention mechanism, allowing for a more focused emphasis on dynamic structural segments that potentially contain attractor structures, thereby enhancing performance. (d) The initialization method proposed for $\bm{A}$ matrix demonstrates marginal yet consistent improvements, underscoring the importance of prior inductive bias for machine learning models. (e) Frequency domain evolution methods significantly reduce temporal noise information and amplify attractor structures. We will further analyze the importance of frequency domain evolution in subsequent analysis.
Chaotic Evolution Strategy.  We further compare various dynamic system evolution strategies mentioned in Section 3.3. From Table 4, it is evident that the frequency-enhanced evolution strategy outperforms the others. Moreover, our proposed efficient implicit evolution method can adaptively explore multi-scale dynamical structure information, avoiding redundant cyclic computations and mitigating overfitting. (b) An inherent characteristic of time series data is significant noise, making it challenging to capture the underlying dynamical structures in the time domain. Direct evolution strategies, whether linear or non-linear neural network-based, do not yield satisfactory results. Moreover, according to the theorem 2 in FiLM [44], the recursion computation for dynamic projection further accumulates noise information. (c) Applying Hopfield networks in the time domain also proves to be unsatisfactory, and even adding more patterns (Key and Value) can have adverse effects. A potential solution is to apply Hopfield networks in the frequency domain instead.
@@ 5 Conclusion and Future Work
LTSF tasks have long been a focal point of research in the machine-learning community. However, mainstream deep learning models currently overlook the crucial aspect that time series data is derived from discretely sampling underlying continuous dynamic systems. Inspired by chaotic theory, our model, Attraos, considers time series as generated by a generalized chaotic dynamic system. By leveraging the invariance of attractors, Attraos enhances predictive performance and provides empirical and theoretical explanations. We aim for Attraos to offer the machine-learning community a fresh perspective and inspire further research on the essence of time series dynamics.
There are still several unresolved issues that require further exploration. For instance: (a) PSR techniques exhibit sensitivity to the numerical values of time series data, and even for the same dataset, different embedding hyperparameters can be obtained. To address this, one potential solution is to utilize implicit neural representations for the phase space coordinates, enabling a trainable embedding process. (b) Due to practical limitations, we have not yet verified our theory on very large-scale chaotic datasets. This aspect will be left for our future research endeavors.
@@ Appendix A Technical Background
A.1 Taken Theorem
In the study of dynamic systems, a delay embedding theorem gives the conditions under which a chaotic dynamic system can be reconstructed from a sequence of observations of the state of that system. The reconstruction preserves the properties of the dynamic system that do not change under smooth coordinate changes (i.e., diffeomorphisms), but it does not preserve the geometric shape of structures in phase space.
Takens’ theorem is the 1981 delay embedding theorem of Floris Takens. It provides the conditions under which a smooth attractor can be reconstructed from the observations made with a generic function. Later results replaced the smooth attractor with a set of arbitrary box counting dimension and the class of generic functions with other classes of functions.
It is the most commonly used method for attractor reconstruction. Delay embedding theorems are simpler to state for discrete-time dynamic systems. The state space of the dynamic system is a $v$-dimensional manifold $\mathcal{M}$. The dynamics is given by a smooth map
$\mathcal{F}:M\rightarrow M$
Assume that the dynamics $\mathcal{F}$ has a strange attractor $\mathcal{A}\subset\mathcal{M}$ with box-counting dimension $d_{A}$. Using ideas from Whitney’s embedding theorem [32], $A$ can be embedded in $m$-dimensional Euclidean space with

### v7

@@ 3.3 System Evolution by Attractors
  • Direct Evolution: We employ each representation $x_{t}/s_{t}\in\mathbb{R}^{D\times N}$ as the feature to partition adjacent points in the dynamical system trajectory into the same attractor using the K-means method. Subsequently, we evolve these points using a local operator $\mathcal{K}^{(i)}$.
  • Frequency-enhanced Evolution: Inspired by neuroscience [6], where attractor structures are amplified in the frequency domain, we first obtain the frequency domain representation of the dynamical structure through Fourier transformation. Considering the dominant modes as attractors, we employ ${\mathcal{K}^{(i)}}$ to drive the system’s evolution in the frequency domain.
  • Hopfield Evolution: Hopfield networks [16] are designed specifically for attractor memory retrieval (see Appendix A.2). In our approach, we utilize a modern version [38] of the Hopfield network for the evolution of dynamical systems, employing cross-attention operations. We treat the trainable attractor library as the Key and Value, while different scales of dynamical structure representations $\{s^{\theta},s^{2\theta},...,s^{2^{L}\theta},x^{2^{L}\theta}\}$ serve as the Query, enabling sequence-to-sequence evolution.
@@
Explicit Evolution. Initially, the finest dynamical representation $x^{\theta}$ is obtained using $\overline{\bm{K}}$, and then expanded to multiple scales ${s^{\theta},s^{2\theta},...,s^{2^{L}\theta},x^{2^{L}\theta}}$. By applying the Fourier Transform, we select $M$ low-frequency components as the primary modes. Each mode undergoes linear evolution $\mathcal{K}^{(i)}=W_{i}\in\mathbb{C}^{N\times N}$, followed by back projection to the original scale.
Implicit Evolution. However, explicit evolution methods inevitably increase the time complexity. To address this issue, inspired by the Blelloch algorithm and hierarchical projection, which both utilize tree-like computational graphs, we propose an implicit evolution method.
A Blelloch scan [3, 42] (Figure 3(d)) defines a binary operator $q_{i}\bullet q_{j}:=(q_{j,a}\odot q_{i,a},~{}q_{j,a}\otimes q_{i,b}+q_{j,b})$ used to compute the linear recurrence $x_{k}=\overline{\bm{A}}x_{k-1}+\overline{\bm{B}}u_{k}$. We take $L=4$ as an example:
@@
The process commences by computing the values of variable $x$ at even positions through an upward sweep. Subsequently, these even position values are employed during a downward sweep to calculate the values at odd positions. The corresponding value of $x$ resides in the right node of $r$. Thus, we can modify the binary operator as $q_{i}\bullet q_{j}:=(q_{j,a}\odot q_{i,a},~{}H_{i}\otimes(q_{j,a}\otimes q_{i,b}+q_{j,b}))$, thereby implicitly integrating hierarchical projection into the scanning operation. This leads to the scales of $x_{i}$:
@@
In this paper, we utilize this indirect and efficient hierarchical projection as the default setting. Ultimately, Attraos projects from the polynomial spectral space back to the phase space by employing another gating projection $\bm{W}_{out}\in\mathbb{R}^{B\times L\times N\times 1}$ (15) to $x\in\mathbb{R}^{B\times L\times D\times N}$, and derives the prediction results using an observation function parameterized by $\bm{W}_{h}\in\mathbb{R}^{LD\times H}$ (flattening the patches).
@@ 4 Experiments
In this section, we commence by conducting a comprehensive performance comparison of Attraos against other state-of-the-art models in seven mainstream LTSF datasets along with two typical chaotic datasets, namely Lorenz96-3d and Air-Convection, followed by ablation experiments pertaining to model architectures. Furthermore, leveraging the properties of chaotic dynamical systems, we explore their extended applications, including experiments on chaotic evolution, representation, reconstruction, and modulation (Appendix E). Finally, we provide the complexity analysis and robustness analysis. For detailed information regarding baseline models, dataset descriptions, experimental settings, and hyper-parameter analysis, please refer to Appendix D.
@@ 4.1 Overall Performance
Model Attraos Mamba4TS S-Mamba RWKV-TS GPT-TS Koopa InvTrm PatchTST DLinear (Ours) (Time Emb.) [47]) [17] [55] [32] [31] [33] [52] Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 0.423 0.420 0.444 0.438 0.459 0.453 0.454 0.446 0.457 0.450 0.450 0.443 0.463 0.454 0.434 0.435 0.462 0.458 ETTh2 0.372 0.399 0.386 0.410 0.381 0.407 0.375 0.402 0.389 0.414 0.397 0.417 0.383 0.407 0.380 0.406 0.564 0.520 ETTm1 0.382 0.391 0.396 0.406 0.399 0.407 0.391 0.403 0.396 0.401 0.395 0.403 0.407 0.412 0.403 0.398 0.403 0.406 ETTm2 0.280 0.324 0.299 0.343 0.289 0.333 0.285 0.330 0.294 0.339 0.281 0.326 0.291 0.335 0.283 0.329 0.345 0.396 Exchange 0.349 0.395 0.364 0.405 0.364 0.407 0.406 0.439 0.371 0.409 0.390 0.424 0.366 0.416 0.383 0.416 0.346 0.416 Crypto 0.187 0.157 0.193 0.162 0.198 0.163 0.190 0.159 0.196 0.164 0.199 0.165 0.196 0.164 0.192 0.161 0.201 0.176 Weather 0.246 0.271 0.258 0.280 0.252 0.277 0.256 0.280 0.279 0.279 0.247 0.273 0.260 0.280 0.258 0.280 0.267 0.319
Mainstream LTSF Datasets. As depicted in Table 1, we can observe that: (a) Attraos consistently exhibits the best performance, closely followed by RWKV-TS and Koopa, which underscores the crucial role of modeling temporal dynamics in LTSF tasks. (b) The models based on state space models (Mamba4TS, S-Mamba) generally outperform the Transformer-based models (PatchTST, InvTrm), indicating the potential superiority of state space models as fundamental frameworks for temporal modeling. (c) The performance of the GPT-TS model, which relies on a pre-trained large language model, is relatively average, suggesting the inherent challenge in directly capturing the dynamics of temporal data using such models. A promising avenue for future research lies in training a dynamical foundational model from scratch on large-scale physical datasets or leveraging pre-training to obtain an attractor tokenizer that is better suited for inputs to the large language model.
@@
Model Attraos Mamba4TS S-Mamba RWKV-TS Koopa InvTrm PatchTST DLinear (Ours) (Time Emb.) [47]) [17] [32] [31] [33] [52] Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Lorenz96-3d 96 0.844 0.684 0.892 0.721 0.925 0.744 0.894 0.722 0.891 0.736 0.963 0.786 0.929 0.756 0.881 0.750 192 0.835 0.662 0.910 0.748 0.917 0.761 0.894 0.744 0.881 0.752 0.944 0.811 0.899 0.714 0.910 0.753 336 0.837 0.681 0.943 0.772 0.968 0.788 0.982 0.823 0.914 0.753 0.997 0.841 0.922 0.787 0.893 0.737 720 0.872 0.739 0.996 0.814 1.135 0.940 1.058 0.921 0.989 0.801 1.129 0.955 0.971 0.828 0.927 0.806 AVG 0.847 0.692 0.935 0.764 0.986 0.808 0.957 0.803 0.919 0.761 1.008 0.848 0.930 0.771 0.903 0.762 Air 96 0.437 0.303 0.451 0.314 0.468 0.329 0.447 0.308 0.443 0.307 0.470 0.337 0.465 0.331 0.441 0.325 192 0.455 0.321 0.472 0.331 0.481 0.340 0.467 0.328 0.451 0.329 0.485 0.349 0.477 0.341 0.460 0.338 336 0.456 0.334 0.468 0.342 0.485 0.351 0.461 0.339 0.468 0.342 0.499 0.363 0.484 0.353 0.461 0.341 720 0.466 0.355 0.492 0.379 0.501 0.386 0.482 0.367 0.488 0.369 0.516 0.401 0.504 0.392 0.474 0.359 AVG 0.454 0.328 0.471 0.342 0.484 0.352 0.464 0.336 0.463 0.337 0.493 0.363 0.483 0.354 0.459 0.341
@@ 4.2 Further Analysis
Ablation Studies.  Next, we turn off each module of Attraos to assess their individual effects. As shown in Table 3, we observe consistent performance decline of Attraos when deleting each module: (a) The removal of Phase Space Reconstruction exhibits the most severe performance degradation, indicating that the process of reconstructing the dynamical structure through PSR forms the foundation for Attraos’ efficient capture of attractor structures. (b) Multi-scale hierarchical projection effectively captures the complex topological structure of the singular attractor, leading to improved performance. However, overfitting may occur in certain prediction lengths. (c) Time-varying $\bm{B}$ and $\bm{W}_{out}$ acts as a gating attention mechanism, allowing for a more focused emphasis on dynamical structure segments that potentially contain attractor structures, thereby enhancing performance. (d) The initialization method proposed for $\bm{A}$ matrix demonstrates marginal yet consistent improvements, underscoring the importance of prior inductive bias for machine learning models. (e) Frequency domain evolution methods significantly reduce temporal noise information and amplify attractor structures. We will further analyze the importance of frequency domain evolution in subsequent analysis.
Chaotic Evolution Strategy.  We further compare various dynamical system evolution strategies mentioned in Section 3.3. From Table 4, it is evident that the frequency-enhanced evolution strategy outperforms the others. Moreover, our proposed efficient implicit evolution method can adaptively explore multi-scale dynamical structure information, avoiding redundant cyclic computations and mitigating overfitting. (b) An inherent characteristic of time series data is significant noise, making it challenging to capture the underlying dynamical structures in the time domain. Direct evolution strategies, whether linear or non-linear neural network-based, do not yield satisfactory results. Moreover, according to the theorem 2 in FiLM [53], the recursion computation for dynamic projection further accumulates noise information. (c) Applying Hopfield networks in the time domain also proves to be unsatisfactory, and even adding more patterns (Key and Value) can have adverse effects. A potential solution is to apply Hopfield networks in the frequency domain instead.
@@
Chaotic Reconstruction.  As illustrated in the left of Figure 5, we visualize the phase space forecasting results of Attraos on the Lorenz96 system. It can be observed that: Although some minor details may be missing due to the sparsity introduced by the frequency domain evolution, Attraos successfully reconstructs the chaotic dynamic structure of Lorenz96. Moreover, modeling time series based on the dynamics structure of the phase space can be viewed as a form of data augmentation, e.g., two-dimensional time figuring [50] or seasonal decomposition [54].
Chaotic Representation & Hyper-parameter Analysis.  In the right of Figure 5, we validated the impact of polynomial dimensions on the model performance. Noteworthy observations include: As noted in substantial literature on SSMs, a polynomial dimension of 256 is generally required to approximate input time series signals with sufficient accuracy. However, we found that the patch operation effectively reduces this threshold in a linear fashion. For instance, in the ETT dataset with a phase space dimension of 4, the required polynomial dimension drops to 256/4 = 64 dimensions.
Figure 5: Left: Chaotic Reconstruction for Lorenz96 system with 720 forecasting step. Right: Hyper-parameter analysis w.r.t. polynomial orders for different model variants w.r.t. patching operation in ETTm2 dataset.
Discussion.  Recently, across various fields of machine learning, an increasing number of works have focused on the underlying physical properties hidden within real-world observational data [20, 21, 26, 51]. By leveraging physical priors, these models achieve significant improvements in generalization, accuracy, and interpretability. We hope Attraos introduces a fresh perspective to the LTSF community and encourages the emergence of physics-guided time series analysis models.
@@ 5 Conclusion and Future Work
LTSF tasks have long been a focal point of research in the machine-learning community. However, mainstream deep learning models currently overlook the crucial aspect that time series data is derived from discretely sampling underlying continuous dynamical systems. Inspired by chaotic theory, our model, Attraos, considers time series as generated by a generalized chaotic dynamical system. By leveraging the invariance of attractors, Attraos enhances predictive performance and provides empirical and theoretical explanations. In the future, we will verify our proposed Attraos on large-scale chaotic datasets and utilize the implicit neural representations for the phase space coordinates, enabling a trainable embedding process to address the instability of PSR techniques.
Acknowledgments and Disclosure of Funding
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
@@ Appendix A Technical Background
A.1 Takens Theorem
Takens’ theorem, introduced by Floris Takens in 1981, provides a framework for reconstructing the dynamics of a chaotic system based on a series of observations. The theorem establishes conditions under which the state space of a dynamical system can be reconstructed from time-delay measurements of a single observable. The reconstructed system retains properties that are invariant under smooth transformations, known as diffeomorphisms.
In the context of discrete-time dynamical systems, Takens’ theorem is commonly applied. Consider a dynamical system whose state space is a $v$-dimensional manifold $\mathcal{M}$, with the evolution of the system governed by a smooth map
$\mathcal{F}:\mathcal{M}\rightarrow\mathcal{M}.$
Assume there exists a strange attractor $\mathcal{A}\subset\mathcal{M}$ with a box-counting dimension $d_{A}$. According to Whitney’s embedding theorem [37], the attractor $\mathcal{A}$ can be embedded into an $m$-dimensional Euclidean space if